Explain what you understand by the apparent magnitude of a star. How is it related
to the brightness of the star? An object A has an apparent magnitude of –5. Another
object B has an apparent magnitude of –10. Calculate the ratio of their brightness
1
Expert's answer
2019-04-15T10:17:13-0400

Apparent magnitude of a star is a measure of its brightness in logarithmic units. By conventional definition, apparent magnitude "m" is related to the observed flux density "F" of light from the star (which is its brightness) by the formula

"m = - 5 \log_{100} \frac{F}{F_0} = - 2.5 \log_{10} \frac{F}{F_0} \, ,"

where "F_0" is the reference flux density corresponding to zero apparent magnitude. Since the flux density (brightness) for a star can be measured in different spectral bands (ultraviolet, visible, infrared etc.), one can speak of the apparent magnitude in the corresponding spectral band. For two objects A and B with apparent magnitudes "m_A" and "m_B", respectively, we have

"m_A - m_B = - 5 \log_{100} \frac{F_A}{F_0} + 5 \log_{100} \frac{F_B}{F_0} = 5 \log_{100} \frac{F_B}{F_A} \, ."

Thus, for "m_A = - 5" and "m_B = - 10", we obtain "5 \log_{100} \left( F_B / F_A \right) = 5", or "\log_{100} \left( F_B / F_A \right) = 1". Hence, the ratio of their brightness is "F_B \/ F_A = 100".